two 10 inch diameter pulleys are 33 inched apart. (center to center of the pulleys) how long a belt is needed to go around them? (inches)
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

two 10 inch diameter pulleys are 33 inched apart. (center to center of the pulleys) how long a belt is needed to go around them? (inches)
-------
Length = 2(33 + semicircle of one pulley)
----
L = 2(33 + pi*5) = 66 + 10pi = 66 + 3.1 = 69.1 inches
-----------------------------
Cheers,
Stan H.
---------------

Answer by MathTherapy(10552)


two 10 inch diameter pulleys are 33 inched apart. (center to center of the pulleys) how long a belt is needed to go around them? (inches)
Correct answer: